OpenStudy (anonymous):

A certain starship can fly 630 miles with the wind in 3 hours and travel the same distance against the wind in 7 hours. What is the speed of the wind?

OpenStudy (dumbcow):

d=r*t Let s be speed of starship, w the speed of wind 630 = 3(s+w) 630 = 7(s-w) solve 1st equation for s s = 210 -w substitute into 2nd equation 630 = 7((210-w)-w) 90 = 210 -2w 2w = 120 w = 60 speed of wind is 60 mph